OpenStudy (anonymous):

Factor completely 16a3b7 + 2a6b4 – 22a4b5

OpenStudy (anonymous):

\[Factor completely 16a ^{3}b ^{7} + 2a ^{6} b ^{4} – 22a ^{4} b ^{5}\]

OpenStudy (anonymous):

\[16 a^3b^7 = 2 \times2\times2\times2 \times a \times a \times a \times b \times b \times b \times b \times b \times b \times b\]\[2a^6b^4 = 2 \times a \times a \times a \times a \times a \times a \times b \times b \times b \times b\]\[22a^4b^5 = 11 \times 2 \times a \times a \times a \times a \times b \times b \times b \times b \times b\] which one are common? O.o

OpenStudy (anonymous):

how many 2, how many a, how many b?

OpenStudy (anonymous):

the answer is 8b^3+b-11ab^2

OpenStudy (anonymous):

16a^3b^7+2a^6b^4–22a^4b^5 not equal 8b^3+b-11ab^2 lol and it's wrong too :c \[2a^3b^4(8b^7 + a^3b-11ab\]common factor

OpenStudy (anonymous):

\[2a^3b^4(8b^3+a^3b−11ab)\]^^^^^^^^^^^^^

OpenStudy (anonymous):

Oh that does make more sense thanks @nphuongsun93